# Concept Based Explanations and Class Contrasting

### v1

@@ Appendix B Generate Visualizations via Gradient Descent
In the main paper we only showed examples where we generate the visualizations by sampling crops from validation images. We can also synthesize them via gradient descent though, similar to (Olah et al., 2017) or (Nguyen et al., 2016). For the following digipath example, we also generate the visualizations for each of the six prototypes via gradient descent in the latent space of a GAN (generative adversarial network). Such visualizations might be more faithful compared to sampling dataset examples. For example in Figure 11 the last third and the last two images show palisading cell edges with different orientations (the first one shows an upper right edge, the next one a lower right edge and the last one a left edge). But in the corresponding dataset visualizations shown in Figure 13 (third row and last two rows), there is no edge orientation visible, but they simply show BCC islands. Meaning, from the dataset visualizations alone, we could not see that the model is actually looking at the edge of the BCC island that has a specific orientation (from only the dataset example we could think that its looking for the islands themselves, or for a right edge instead of a left edge).
Appendix C Basal Cell Carcinoma vs Trichoblastoma
@@
Basal Cell Carcinoma (BCC) as well as Trichoblastoma look clinically very similar to each other (Patel et al., 2020). Human experts as well as the AI model can separate them, and in this section we investigate why the model predicts Basal Cell Carcinoma over Trichoblastoma. First we can look at the main concepts the model uses to predict BCC shown in Figure 11 and compare them to the main concepts the model uses to predict Trichoblastoma shown in Figure 10. The problem is that the concepts look very similar (which makes sense since the two diseases look clinically similar), there is no single concept that is present for one class but not the other. For example the first concept from the left for Trichoblastoma looks similar compared to the first from the left for BCC. The second one for Trichoblastoma looks similar to the fourth for BCC and the third looks similar to the second. And the last three concepts for Trichoblastoma show palisading cell edges with different orientations, similar to the third and last two concepts for BCC. So those visualizations are not enough to spot a difference between Trichoblastoma and BCC.
@@
In the next step we look at the visualizations pro BCC we get from contrasting the two classes. Figure 12 shows the 6 main concepts that are pro BCC vs Trichoblastoma. The second, third and fourth concept look differently from the 6 main concepts pro Trichoblastoma (so here we can now see a difference). For interpreting those 3 concepts we can look at dataset examples, shown in Figure 14. The fourth concept shows white areas inside tumor, which could resemble cleavage (in BCC the tumor is often separated from the surrounding tissue by white area). The third concept shows epidermis, which is also interesting since according to (Patel et al., 2020) BCC is sometimes attached to the epidermis, whereas Trichoblastoma never is.
For the quantitative check, the average maximum prediction for BCC in Trichoblastoma images is 0.013 and the average maximum prediction for Trichoblastoma is 0.889. So at this point the model is very sure that the images are Trichoblastoma and not BCC. When masking 20% of the input pixels that lie furthest on the Trichoblastoma side of the hyperplane between Trichoblastoma and BCC, then the maximum average prediction for BCC goes up to 0.408, which is an increase by a factor of 31. And the average maximum prediction for Trichoblastoma goes down to 0.467, so now the model is no longer sure whether the images belong to BCC or Trichoblastoma. We mask the input pixels in steps of 10%, and masking more or less than those 20% reduces the prediction for BCC. Something to note, is that between BCC and Trichoblastoma this masking only works one way, from Trichoblastoma to BCC. Doing it the other way around (using BCC images and mask the most important features for BCC vs Trichoblastoma) reduces the predictions for both BCC as well as Trichoblastoma.
@@
Shifting the activations in the latent space in the direction of the hyperplane normal works both ways. Here we reach an average maximum prediction for BCC of 0.721 (at that point the prediction for Trichoblastoma is 0.481). And the other way around, we reach an average maximum prediction for Trichoblastoma of 0.673 and at that point the prediction for BCC is 0.017.
Appendix D Examples for Single Class for the ResNet50
@@
Figure 15: Visualization for the 6 NMF components for fireboat, using layer4.2. The model predicts 1.0 for fireboat for the six prototypes (passing the first column into the model)
@@
Figure 16: Visualization for the 6 NMF components for fireboat, using layer4.2 and excluding images where the model predicts fireboat. The model predicts 0.98 for fireboat for the six prototypes
@@
Figure 17: Visualization for the 6 NMF components for gazelle, using layer3.5. The model predicts 1.0 for gazelle for the six prototypes
@@
Figure 18: Visualization for the 6 NMF components for gazelle, using layer3.5 and excludig images where the model predicts gazelle. The model predicts 0.82 for gazelle for the six prototypes
@@
Figure 19: Visualization for the 6 NMF components for gazelle, using layer4.2. The model predicts 0.99 for gazelle for the six prototypes
@@
Figure 20: Visualization for the 6 NMF components for gazelle, using layer4.2 and excluding images where the model predicts gazelle. Now the model predicts only 0.33 for gazelle, and 0.67 for impala. This would be a failure case, because gazelle is not the majority class

### v2

@@ Appendix E Generate Visualizations via Gradient Descent
In the main paper we only showed examples where we generate the visualizations by sampling crops from validation images. We can also synthesize them via gradient descent though, similar to Olah et al. [2017] or Nguyen et al. [2016]. For the following digital pathology example, we also generate the visualizations for each of the six concepts via gradient descent in the latent space of a GAN (generative adversarial network). Such visualizations might be more faithful compared to sampling dataset examples. For example in Figure 15 the last third and the last two images show palisading cell edges with different orientations (the first one shows an upper right edge, the next one a lower right edge and the last one a left edge). But in the corresponding dataset visualizations shown in Figure 17 (third row and last two rows), there is no edge orientation visible, they simply show BCC islands. Meaning, from the dataset visualizations alone, we could not see that the model is actually looking at the edge of the BCC island that has a specific orientation (from only the dataset example we could think that its looking for the islands themselves, or for a right edge instead of a left edge).
Appendix F Basal Cell Carcinoma vs Trichoblastoma
@@
Basal Cell Carcinoma (BCC) as well as Trichoblastoma look clinically very similar to each other Patel et al. [2020]. Human experts as well as the AI model can separate them, and in this section we investigate why the model predicts Basal Cell Carcinoma over Trichoblastoma. First we can look at the main concepts the model uses to predict BCC shown in Figure 15 and compare them to the main concepts the model uses to predict Trichoblastoma shown in Figure 14. The problem is that the concepts look very similar (which makes sense since the two diseases look clinically similar), there is no single concept that is present for one class but not the other. For example the first concept from the left for Trichoblastoma looks similar compared to the first from the left for BCC. The second one for Trichoblastoma looks similar to the fourth for BCC and the third looks similar to the second. And the last three concepts for Trichoblastoma show palisading cell edges with different orientations, similar to the third and last two concepts for BCC. So those visualizations are not enough to spot a difference between Trichoblastoma and BCC.
@@
In the next step we look at the visualizations pro BCC we get from contrasting the two classes. Figure 16 shows the 6 main concepts that are pro BCC vs Trichoblastoma. The second, third and fourth concept look differently from the 6 main concepts pro Trichoblastoma (so here we can now see a difference). For interpreting those 3 concepts we can look at dataset examples, shown in Figure 18. The fourth concept shows white areas inside tumor, which could resemble cleavage (in BCC the tumor is often separated from the surrounding tissue by white area). The third concept shows epidermis, which is also interesting since according to Patel et al. [2020] BCC is sometimes attached to the epidermis, whereas Trichoblastoma never is.
@@
For the shifting test, shifting the activations in the latent space in the direction of the hyperplane normal works both ways. Here we reach an average maximum prediction for BCC of 0.721 (at that point the prediction for Trichoblastoma is 0.481). And the other way around, we reach an average maximum prediction for Trichoblastoma of 0.673 and at that point the prediction for BCC is 0.017.
Appendix G Examples ResNet50
Figure 19: Concept based explanation for the class ’computer keyboard, keypad’. The second concept shows cats (in front of a monitor) which signifies a bias in the model/data, since the presence of cats should not be used to predict the presence of a computer keyboard
Figure 20: Concept based explanation for the class ’horizontal bar, high bar’
@@ Appendix G Examples ResNet50
Figure 21: Visualization for the 4 NMF components for orange (fruit), using layer4.2. The model predicts 1.0 for orange for the four concepts (passing the first column into the model)
@@
Figure 22: Visualization for the 4 NMF components for orange, using layer4.2 and excluding images where the model predicts orange. The model predicts 0.97 for orange for the four concepts
In this section we include some more examples for the ResNet50 model. For example, Figure 19 shows a concept based explanation for why the model predicts ’computer keyboard, keypad’. Here the second concept shows cats in front of a monitor, meaning the model utilizes the presence of cats when predicting the class ’computer keyboard, keypad’. That can be problematic, because if is something is not a computer keyboard it does not become one if a cat is next to it and a computer keyboard is still one even if there is no cat next to it. Or Figure 20 shows a concept based explanation for the class ’horizontal bar, high bar’, but the first concept does not show any bar at all, only gymnasts. The second one shows hands on a bar and the third one shows small children hanging from a bar. Meaning the model will likely have problems correctly classifying a horizontal bar if there is no human on it.
Figure 22 shows the four main concepts (excluding patches where the model predicts orange) for orange (fruit) for layer4.2 of the ResNet50 model from pytorch model zoo. Even taking the bottom two concepts (the edge of an orange ping pong ball and a bunch of fruits) combined gets predicted as orange with a confidence of 1.0. That points towards a problem in the model for this class (bunch of fruits plus edge of an orange ball gets predicted as orange). Also none of the four concepts actually shows orange as a plant, but only fruits or slices, which might also points towards a bias in the model.
@@
Figure 23: Visualization for the 4 NMF components for fireboat, using layer4.2. The model predicts 1.0 for fireboat for the four concepts (passing the first column into the model)
@@
Figure 24: Visualization for the 4 NMF components for fireboat, using layer4.2 and excluding images where the model predicts fireboat. The model still predicts 1.0 for fireboat for the four concepts
@@
Figure 25: Visualization for the 4 NMF components for gazelle, using layer4.2. The model predicts 1.0 for gazelle for the four concepts
@@
Figure 26: Visualization for the 4 NMF components for gazelle, using layer4.2 and excluding images where the model predicts gazelle. The model only predicts 0.05 for gazelle, with impala being the majority class with 0.95